9-4{-2(4-8)-5(3-1)}
Answer by Alan3354(69443)

9-4{-2(4-8)-5(3-1)}
= 9-4{-2(-4)-5(2)}
= 9 - 4{8 - 10}
= 9 - 4(-2)
= 9 + 8
= 17